# Splittings of One-Ended Groups with One-Ended Halfspaces

### v3

@@ 1 Introduction
The idea of an “artificial” splitting can be used to transform a non-trivial splitting over a subgroup $C$ into a non-trivial splitting over a larger group. We use artificial splittings to produce an example (see Example 7.2) of a finitely presented one-ended simply connected at infinity group $G$ and a non-trivial splitting of $G$ over an infinite-ended group. So, it is important to know when halfspaces are one-ended.
Perhaps the most important application of the notion of simple connectivity at infinity is to manifolds. The notion of simple connectivity at infinity was popularized by J. Stallings [Sta62], J. Munkres [Mun60], E. Luff [Luf67] and L. Siebenmann [Sie68] in the 1960’s as a way of deciding when a contractible open $n$-manifold is homeomorphic to $\mathbb{R}^{n}$ for $n\geq 5$. Following the work of Freedman in dimension 4 (see Corollary1.2 of [Fre82]) and Perelman in dimension 3 (see L. Husch and T. Price’s paper [HP70] and the addendum [HP71]), these results were extended to all dimensions $n>2$. So for contractible open topological, PL and differential manifolds $M$ of dimension $n>2$ we have that $M$ is homeomorphic to $\mathbb{R}^{n}$ if and only if $M$ is simply connected at infinity. (Note that in dimension 4 there are exotic PL and differential structures on $\mathbb{R}^{4}$ see S. Donaldson’s paper [Don83].) M. Davis [Dav83] produced closed manifolds with infinite fundamental group in dimensions $n\geq 4$ whose universal covers were contractible, and he proved these universal covers are not simply connected at infinity (and hence not homeomorphic to $\mathbb{R}^{n}$). The fundamental group $G$ of a Davis manifold is a subgroup of finite index of a right angled Coxeter group $W$ and so many “visual” splittings of $W$ (and $G$) are easily obtained. However, we have a restriction on splittings of such groups by the following corollary of Corollary 1.5.
@@
§7 begins with the notion of an artificial splitting of a group. We prove a lemma that describes how to alter a given non-trivial amalgamated product to obtain a non-trivial one with larger edge and vertex groups. Example 7.2 begins with a finitely presented group $G$ that is simply connected at infinity and splits as an amalgamated product with simply connected at infinity vertex groups and a one-ended edge group. An artificial splitting is constructed for $G$ with an infinite-ended vertex group and an infinite-ended edge group. So simply connected at infinity groups can be split non-trivially over infinite-ended groups. Corollary 1.5 implies a halfspace for this last splitting is not one-ended (so halfspaces for one-ended simply connected at infinity groups need not always be one-ended).
@@ 2 Group splittings and their halfspaces
A similar definition can be made for HNN extensions, as follows. Suppose $G=A*_{C}$ is an HNN extension, with $A$ and $C$ finitely generated. Let $S_{A}$ be a finite generating set for $A$, and let $t$ be the stable letter of the HNN extension. Let $\operatorname{Cay}(G,S)$ denote the Cayley graph of $G$ with respect to $S=S+A\cup\{t\}$. Define the halfspace $\mathfrak{h}^{+}$ to be the induced subgraph of $\operatorname{Cay}(G,S)$ whose vertices are elements of $G$ with normal forms that do not start with an element of $C$ followed by a negative power of $t$. Define the halfspace $\mathfrak{h}^{-}$ to be the induced subgraph of $\operatorname{Cay}(G,S)$ whose vertices are either elements of $C$ or elements of $G$ with normal forms that do start with an element of $C$ followed by a negative power of $t$. We refer to $\mathfrak{h}^{+},\mathfrak{h}^{-}$ and all of their $G$-translates as halfspaces of the splitting $G=A*_{C}$. Note that these halfspaces are connected subgraphs of $\operatorname{Cay}(G,S)$, and $\mathfrak{h}^{+}\cap\mathfrak{h}^{-}$ is the induced subgraph of $C$.
@@
Let $G\curvearrowright T$ be a splitting of a finitely generated group $G$ with minimal action and finitely generated edge stabilizers. Fix an orbit map $f:G\to VT$, and a Cayley graph $\operatorname{Cay}(G,S)$ of $G$ with respect to a finite generating set $S$. Define a halfspace of $G\curvearrowright T$ (with respect to $S$) to be an induced subgraph of $\operatorname{Cay}(G,S)$ with vertex set a preimage $f^{-1}(\mathfrak{h})$ of a halfspace $\mathfrak{h}\in\mathcal{H}(T)$. If $\mathfrak{h}$ is bounded by the edge midpoint $\hat{e}$ then we say that $f^{-1}(\mathfrak{h})$ is a halfspace of $G\curvearrowright T$ associated to $e$. For an appropriate choice of $S$, all the halfspaces of $G\curvearrowright T$ will be connected subgraphs of $\operatorname{Cay}(G,S)$. (If $S_{R}$ is the $f$-preimage of the $R$-ball in $T$ centered at $f(1)$, then one can show that all halfspaces of $G\curvearrowright T$ with respect to $S_{R}$ will be connected for large enough $R$ (extending the definition of halfspace to infinite generating sets); and if a finite generating set $S$ for $G$ contains a generating set for $G_{f(1)}$ and satisfies $S_{R}\subseteq G_{f(1)}S$, then one can show that all halfspaces of $G\curvearrowright T$ with respect to $S$ will be connected.)
@@
We now prove that these three definitions of halfspace are well-defined and equivalent up to quasi-isometry. This is not as straightforward to prove as one might imagine, since the inclusion of a halfspace $\mathfrak{h}\xhookrightarrow{}\operatorname{Cay}(G,S)$ (say for the second definition) is not necessarily a quasi-isometric embedding.
@@
The three definitions for halfspaces given above are coarsely well-defined and coarsely equivalent. More precisely, if we fix a splitting $G\curvearrowright T$ and a halfspace $\mathfrak{h}$ of $T$, then the halfspace of $G\curvearrowright T$ corresponding to $\mathfrak{h}$ remains the same up to quasi-isometry when we change the choice of finite generating set for $G$ and when we switch between the three definitions of halfspace (and when we change the orbit map $f$ from the second definition and the graph $X$ from the third definition). In particular, the notion of the halfspaces of a splitting $G\curvearrowright T$ being one-ended is well-defined.
@@
We begin by showing that the first definition is a special case of the second. Let $G=A*_{C}B$ be a non-trivial amalgam with $A,B$ and $C$ finitely generated, and let $G\curvearrowright T$ be the corresponding Bass–Serre tree. Let $v_{A},v_{B}\in VT$ be the vertices stabilized by $A$ and $B$ respectively, and let $e$ be the edge joining $v_{A}$ and $v_{B}$. Let $\mathfrak{h}\subseteq T$ be the halfspace bounded by $\hat{e}$ that contains $v_{A}$. Let $f:G\to VT$ be the orbit map given by $g\mapsto gv_{A}$ and let $S=S_{A}\cup S_{B}$ be the finite generating set for $G$ from the first definition. Then it is not hard to show that the halfspace $\mathfrak{h}_{A}$ from the first definition is exactly the same as the halfspace from the second definition with vertex set $f^{-1}(\mathfrak{h})$ (with respect to generating set $S$). Similar arguments for $\mathfrak{h}_{B}$ and for the $G$-translates of $\mathfrak{h}_{A}$ and $\mathfrak{h}_{B}$ show that the set of halfspaces from the first definition is the same as the set of halfspaces from the second definition. A similar argument works for HNN extensions.
Next, we show the equivalence of the second and third definitions. Let $e\in ET$ and let $\mathfrak{h}\in\mathcal{H}(T)$ be one of the halfspaces bounded by $\hat{e}$. We have an orbit map $f:G\to T$ from the second definition and a tree of spaces $p:X\to T$ from the third definition. Let $i:G\to VX$ be an orbit map for the action of $G$ on $X$ such that the following diagram commutes.
${G}$${X}$${T}$$\scriptstyle{f}$$\scriptstyle{i}$$\scriptstyle{p}$ (1)
The second definition provides a halfspace $\mathfrak{h}_{1}\subseteq\operatorname{Cay}(G,S)$ (with respect to a suitable finite generating set $S$) with vertex set $V\mathfrak{h}_{1}=f^{-1}(\mathfrak{h})$, while the third definition provides a halfspace $\mathfrak{h}_{2}:=p^{-1}(\mathfrak{h})\subseteq X$. Denote the metrics on $\operatorname{Cay}(G,S)$ and $X$ by $d_{S}$ and $d_{X}$ respectively, and denote the induced path metrics on $\mathfrak{h}_{1}$ and $\mathfrak{h}_{2}$ by $d_{\mathfrak{h}_{1}}$ and $d_{\mathfrak{h}_{2}}$ respectively. Observe that $i(V\mathfrak{h}_{1})\subseteq V\mathfrak{h}_{2}$. We claim that the map $i:(V\mathfrak{h}_{1},d_{\mathfrak{h}_{1}})\to(V\mathfrak{h}_{2},d_{\mathfrak{h}_{2}})$ is a quasi-isometry.
The Milnor–Svarc Lemma implies that $i:(G,d_{S})\to(X,d_{X})$ is a quasi-isometry. Say $C\geq 1$ is such that
@@
Let $\hat{\mathfrak{h}}_{2}:=p^{-1}(\hat{e})\subseteq\mathfrak{h}_{2}$. Since $G_{e}$ acts cocompactly on the $4C$-neighborhood of $\hat{\mathfrak{h}}_{2}$, preserving $\mathfrak{h}_{2}$, there is a constant $L$ such that, for any $v,v^{\prime}\in V\mathfrak{h}_{2}$, if $d_{X}(v,v^{\prime})\leq 2C$ and $d_{X}(v,\hat{\mathfrak{h}}_{2})\leq 2C$ then $d_{\mathfrak{h}_{2}}(v,v^{\prime})\leq L$. We may also assume $L\geq 2C$. Now consider adjacent vertices $g,g^{\prime}\in V\mathfrak{h}_{1}$. By (2), we have $d_{X}(i(g),i(g^{\prime}))\leq Cd_{S}(g,g^{\prime})+C=2C$. Note that any path leaving $\mathfrak{h}_{2}$ must cross $\hat{\mathfrak{h}}_{2}$, so if $d_{X}(i(g),\hat{\mathfrak{h}}_{2})\geq 2C$ then $i(g^{\prime})\in\mathfrak{h}_{2}$ and $d_{\mathfrak{h}_{2}}(i(g),i(g^{\prime}))=d_{X}(i(g),i(g^{\prime}))\leq 2C$. Otherwise, if $d_{X}(i(g),\hat{\mathfrak{h}}_{2})<2C$ then $d_{\mathfrak{h}_{2}}(i(g),i(g^{\prime}))\leq L$ by definition of $L$. In either case, we have $d_{\mathfrak{h}_{2}}(i(g),i(g^{\prime}))\leq L$. By mapping paths in $\mathfrak{h}_{1}$ into $X$, we deduce that

### v4

@@ 1 Introduction
We remark that, in the finitely presented case, Corollary 1.2 can alternatively be deduced from a result of Papasoglu [Pap05, Definition 1.4 and Lemma 1.8] (note that Lemma 1.8 in [Pap05] is not quite stated correctly: the group $G$ should be one-ended).
@@
By folding, one can transform a non-trivial splitting over a subgroup $C$ into a non-trivial splitting over a larger subgroup. We use this idea to produce an example (see Example 7.2) of a finitely presented one-ended simply connected at infinity group $G$ and a non-trivial splitting of $G$ over an infinite-ended group. So, it is important to know when halfspaces are one-ended.
Perhaps the most important application of the notion of simple connectivity at infinity is to manifolds. The notion of simple connectivity at infinity was popularized by J. Stallings [Sta62], J. Munkres [Mun60], E. Luff [Luf67] and L. Siebenmann [Sie68] in the 1960’s as a way of deciding when a contractible open $n$-manifold is homeomorphic to $\mathbb{R}^{n}$ for $n\geq 5$. Following the work of Freedman in dimension 4 (see Corollary1.2 of [Fre82]) and Perelman in dimension 3 (see L. Husch and T. Price’s paper [HP70] and the addendum [HP71]), these results were extended to all dimensions $n>2$. So for contractible open topological, PL and differential manifolds $M$ of dimension $n>2$ we have that $M$ is homeomorphic to $\mathbb{R}^{n}$ if and only if $M$ is simply connected at infinity. (Note that in dimension 4 there are exotic PL and differential structures on $\mathbb{R}^{4}$ - see S. Donaldson’s paper [Don83].) M. Davis [Dav83] produced closed manifolds with infinite fundamental group in dimensions $n\geq 4$ whose universal covers were contractible, and he proved these universal covers are not simply connected at infinity (and hence not homeomorphic to $\mathbb{R}^{n}$). The fundamental group $G$ of a Davis manifold is a subgroup of finite index of a right angled Coxeter group $W$ and so many “visual” splittings of $W$ (and $G$) are easily obtained. However, we have a restriction on splittings of such groups by the following corollary of Corollary 1.5.
@@
§7 begins with a lemma that describes how to alter a given non-trivial amalgamated product to obtain a non-trivial one with larger edge and vertex groups. Example 7.2 begins with a finitely presented group $G$ that is simply connected at infinity and splits as an amalgamated product with simply connected at infinity vertex groups and a one-ended edge group. A second splitting is constructed for $G$ with an infinite-ended vertex group and an infinite-ended edge group. So simply connected at infinity groups can be split non-trivially over infinite-ended groups. Corollary 1.5 implies that a halfspace for this last splitting is not one-ended (so halfspaces for one-ended simply connected at infinity groups need not always be one-ended).
@@
Acknowledgements:  We thank Alex Margolis for suggesting Example 2.8. And we are grateful for the referee’s careful reading of our paper and their helpful comments and corrections.
@@ 2 Group splittings and their halfspaces
A similar definition can be made for HNN extensions, as follows. Suppose $G=A*_{C}$ is an HNN extension, with $A$ and $C$ finitely generated. Let $S_{A}$ be a finite generating set for $A$, and let $t$ be the stable letter of the HNN extension. Let $\operatorname{Cay}(G,S)$ denote the Cayley graph of $G$ with respect to $S=S_{A}\cup\{t\}$. Define the halfspace $\mathfrak{h}^{+}$ to be the induced subgraph of $\operatorname{Cay}(G,S)$ whose vertices are elements of $G$ with normal forms that do not start with an element of $C$ followed by a negative power of $t$. Define the halfspace $\mathfrak{h}^{-}$ to be the induced subgraph of $\operatorname{Cay}(G,S)$ whose vertices are either elements of $C$ or elements of $G$ with normal forms that do start with an element of $C$ followed by a negative power of $t$. We refer to $\mathfrak{h}^{+},\mathfrak{h}^{-}$ and all of their $G$-translates as halfspaces of the splitting $G=A*_{C}$. Note that these halfspaces are connected subgraphs of $\operatorname{Cay}(G,S)$, and $\mathfrak{h}^{+}\cap\mathfrak{h}^{-}$ is the induced subgraph of $C$.
@@
Let $G\curvearrowright T$ be a splitting of a finitely generated group $G$ with minimal action and finitely generated edge stabilizers. Fix an orbit map $f:G\to VT$. Let $S$ be a finite generating set for $G$ such that the induced subgraph of $f^{-1}(\mathfrak{h})$ in $\operatorname{Cay}(G,S)$ is connected for every halfspace $\mathfrak{h}\in\mathcal{H}(T)$ – such $S$ exists by Lemma 2.4 below. These induced subgraphs are called the halfspaces of $G\curvearrowright T$ (with respect to $S$). If $\mathfrak{h}$ is bounded by the edge midpoint $\hat{e}$ then we say that the halfspace of $G\curvearrowright T$ induced by $f^{-1}(\mathfrak{h})$ is a halfspace associated to $e$.
Lemma 2.4.
Let $G\curvearrowright T$ be a splitting of a finitely generated group $G$ with minimal action and finitely generated edge stabilizers. Fix an orbit map $f:G\to VT$. There exists a finite generating set $S$ for $G$ such that the induced subgraph of $f^{-1}(\mathfrak{h})$ in $\operatorname{Cay}(G,S)$ is connected for every halfspace $\mathfrak{h}\in\mathcal{H}(T)$.
Proof.
Let $\mathfrak{h}\in\mathcal{H}(T)$ be a halfspace associated to an edge $e\in ET$. We will find a finite generating set $S$ for $G$ such that the induced subgraph of $f^{-1}(\mathfrak{h})$ in $\operatorname{Cay}(G,S)$ is connected. The lemma then follows by repeating the argument for a finite set of $G$-orbit representatives in $\mathcal{H}(T)$, and taking the union of the corresponding finite generating sets for $G$.
Let $S_{0}$ be a finite symmetric generating set for $G$, and let
$R=\max\{d(f(1),f(s))\mid s\in S_{0}\}=\max\{d(f(g),f(gs))\mid g\in G,s\in S_{0}\}.$ (1)
Let $\mathfrak{h}^{*}\in\mathcal{H}(T)$ be the halfspace which is complementary to $\mathfrak{h}$, and define
$X_{e}=\{(g,gs)\in G^{2}\mid f(g)\in\mathfrak{h},f(gs)\in\mathfrak{h}^{*},s\in S_{0}\}.$ (2)
Since the orbit map $f$ is $G$-equivariant, the left action of $G_{e}$ on $G^{2}$ stabilizes $X_{e}$. We claim that $X_{e}$ contains only finitely many $G_{e}$-orbits. Indeed, suppose $s\in S_{0}$ and $(g_{1},g_{1}s),(g_{2},g_{2}s)\in X_{e}$ with $d(f(g_{1}),e)=d(f(g_{2}),e)$. The element $g_{2}g_{1}^{-1}$ maps the geodesic in $T$ from $f(g_{1})$ to $f(g_{1}s)$ onto the geodesic from $f(g_{2})$ to $f(g_{2}s)$. Since $d(f(g_{1}),e)=d(f(g_{2}),e)$, it must be that $g_{2}g_{1}^{-1}\in G_{e}$, so $(g_{1},g_{1}s),(g_{2},g_{2}s)\in X_{e}$ are in the same $G_{e}$-orbit. For $(g,gs)\in X_{e}$, there are only finitely many possibilities for the choice of $s\in S_{0}$, and the distance $d(f(g),e)$ is bounded by $R$ from (1) because $e$ lies on the geodesic in $T$ from $f(g)$ to $f(gs)$. This proves the claim that $X_{e}$ contains only finitely many $G_{e}$-orbits.
Now let $\{(g_{i},g_{i}s_{i})\}_{i=1}^{n}$ be a finite set of $G_{e}$-orbit representatives in $X_{e}$. Let $S_{e}$ be a finite generating set for $G_{e}$, and let
$S_{e}^{\prime}=\bigcup_{i=1}^{n}g_{i}^{-1}S_{e}g_{i}\quad\text{and}\quad S_{e}^{\#}=\{g_{i}^{-1}g_{j}\mid i,j\in\{1,\dots,n\}\}.$ (3)
We claim that, for each $i=1,\dots,n$ and each $h\in G_{e}$, there is a path from $g_{i}$ to $hg_{i}$ in $\operatorname{Cay}(G,S_{0}\cup S_{e}^{\prime})$ that stays in $f^{-1}(\mathfrak{h})$. Indeed, write $h=s_{1}s_{2}\cdots s_{k}$ with $s_{j}\in S_{e}$. Then the desired path is
$(g_{i},g_{i}(g_{i}^{-1}s_{1}g_{i}),g_{i}(g_{i}^{-1}s_{1}g_{i})(g_{i}^{-1}s_{2}g_{i}),\dots,g_{i}(g_{i}^{-1}s_{1}g_{i})(g_{i}^{-1}s_{2}g_{i})\cdots(g_{i}^{-1}s_{k}g_{i}))$
$=$ $(g_{i},s_{1}g_{i},s_{1}s_{2}g_{i},\dots,hg_{i}).$
Let $S=S_{0}\cup S_{e}^{\prime}\cup S_{e}^{\#}$. It follows from the previous claim that, for any $(g,gs),(g^{\prime},g^{\prime}s^{\prime})\in X_{e}$, there is a path from $g$ to $g^{\prime}$ in $\operatorname{Cay}(G,S)$ that stays in $f^{-1}(\mathfrak{h})$.
We now complete the proof of the lemma by showing that the induced subgraph of $f^{-1}(\mathfrak{h})$ in $\operatorname{Cay}(G,S)$ connected. Indeed, let $g,g^{\prime}\in f^{-1}(\mathfrak{h})$. Take a path $\gamma$ in $\operatorname{Cay}(G,S_{0})$ from $g$ to $g^{\prime}$, and consider it as a path in $\operatorname{Cay}(G,S)$. Each maximal subpath of $\gamma$ that lies in $f^{-1}(\mathfrak{h}^{*})$ must be preceded by an edge $(g^{\prime\prime},g^{\prime\prime}s^{\prime\prime})\in f^{-1}(\mathfrak{h})\times f^{-1}(\mathfrak{h}^{*})$ and immediately followed by an edge $(g^{\prime\prime\prime},g^{\prime\prime\prime}s^{\prime\prime\prime})\in f^{-1}(\mathfrak{h}^{*})\times f^{-1}(\mathfrak{h})$ (with $s^{\prime\prime},s^{\prime\prime\prime}\in S_{0}$). Since $S_{0}$ is symmetric, we have $(g^{\prime\prime},g^{\prime\prime}s^{\prime\prime}),(g^{\prime\prime\prime}s^{\prime\prime\prime},g^{\prime\prime\prime})\in X_{e}$. Hence it follows from the arguments above that the subpath of $\gamma$ from $g^{\prime\prime}$ to $g^{\prime\prime\prime}s^{\prime\prime\prime}$ can be replaced by a path in $f^{-1}(\mathfrak{h})$. Doing this for every maximal subpath of $\gamma$ that lies in $f^{-1}(\mathfrak{h}^{*})$, we obtain a path from $g$ to $g^{\prime}$ in $f^{-1}(\mathfrak{h})$, as required. ∎
@@
In the following three lemmas we prove that these three definitions of halfspace are well-defined and equivalent up to quasi-isometry. This is not as straightforward to prove as one might imagine, since the inclusion of a halfspace $\mathfrak{h}\xhookrightarrow{}\operatorname{Cay}(G,S)$ (say for the second definition) is not necessarily a quasi-isometric embedding (see Example 2.8).
@@
Let $G\curvearrowright T$ be a splitting of a finitely generated group $G$ with minimal action and finitely generated edge stabilizers. Fix a finite generating set $S$ and an orbit map $f:G\to VT$. Let $p:X\to T$ be a tree of spaces as in Definition 2. Fix an edge $e\in ET$ and let $\mathfrak{h}\in\mathcal{H}(T)$ be one of the halfspaces associated to $e$. Suppose that the induced subgraph $\mathfrak{h}_{1}$ of $f^{-1}(\mathfrak{h})$ in $\operatorname{Cay}(G,S)$ is connected (but make no assumption regarding the other halfspaces of $T$). Let $\mathfrak{h}_{2}=p^{-1}(\mathfrak{h})\subseteq X$ be the halfspace corresponding to $\mathfrak{h}$ from Definition 2. Let $i:G\to VX$ be an orbit map for the action of $G$ on $X$ such that the following diagram commutes.
[…]
Then the map $i:(V\mathfrak{h}_{1},d_{\mathfrak{h}_{1}})\to(V\mathfrak{h}_{2},d_{\mathfrak{h}_{2}})$ is a quasi-isometry, where $d_{\mathfrak{h}_{1}},d_{\mathfrak{h}_{2}}$ denote the induced path metrics on $\mathfrak{h}_{1},\mathfrak{h}_{2}$ respectively.
@@
Denote the metrics on $\operatorname{Cay}(G,S)$ and $X$ by $d_{S}$ and $d_{X}$ respectively. The Milnor–Svarc Lemma implies that $i:(G,d_{S})\to(X,d_{X})$ is a quasi-isometry. Say $C\geq 1$ is such that
@@
Let $\hat{\mathfrak{h}}_{2}:=p^{-1}(\hat{e})\subseteq\mathfrak{h}_{2}$. By Definition 2, $G$ acts cocompactly on $X$, hence $G_{e}$ acts cocompactly on $\hat{\mathfrak{h}}_{2}$. As $X$ is locally finite, $G_{e}$ also acts cocompactly on the $4C$-neighborhood of $\hat{\mathfrak{h}}_{2}$. It follows that there is a constant $L$ such that, for any $v,v^{\prime}\in V\mathfrak{h}_{2}$, if $d_{X}(v,v^{\prime})\leq 2C$ and $d_{X}(v,\hat{\mathfrak{h}}_{2})\leq 2C$ then $d_{\mathfrak{h}_{2}}(v,v^{\prime})\leq L$. We may also assume $L\geq 2C$. Now consider adjacent vertices $g,g^{\prime}\in V\mathfrak{h}_{1}$. By (5), we have $d_{X}(i(g),i(g^{\prime}))\leq Cd_{S}(g,g^{\prime})+C=2C$. Note that any path leaving $\mathfrak{h}_{2}$ must cross $\hat{\mathfrak{h}}_{2}$, so if $d_{X}(i(g),\hat{\mathfrak{h}}_{2})\geq 2C$ then $i(g^{\prime})\in\mathfrak{h}_{2}$ and $d_{\mathfrak{h}_{2}}(i(g),i(g^{\prime}))=d_{X}(i(g),i(g^{\prime}))\leq 2C$. Otherwise, if $d_{X}(i(g),\hat{\mathfrak{h}}_{2})<2C$ then $d_{\mathfrak{h}_{2}}(i(g),i(g^{\prime}))\leq L$ by definition of $L$. In either case, we have $d_{\mathfrak{h}_{2}}(i(g),i(g^{\prime}))\leq L$. By mapping paths in $\mathfrak{h}_{1}$ into $X$, we deduce that